OpenStudy (anonymous):

2 coordinate algebra questions!

OpenStudy (anonymous):

1. Which points are on the graph of g(x)=(1/5)^x? Choose all answers that are correct. (-2,1/25) (-1,5) (1,0) (3,1/125)

OpenStudy (anonymous):

@TheSmartOne Can you please help?

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/anndmguobm

OpenStudy (anonymous):

I still don't get it, sorry.

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/eomlzedeuk

TheSmartOne (thesmartone):

Zoom into each point... If it is on the line it is the answer.... There is only one answer for this quesion...

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

what is .2^1

TheSmartOne (thesmartone):

x^1=x

OpenStudy (anonymous):

2^1=2

TheSmartOne (thesmartone):

yes but it is.2 So that is not the correct answer...

TheSmartOne (thesmartone):

Well I gtg...

TheSmartOne (thesmartone):

Its B...

TheSmartOne (thesmartone):

If you look at this grph https://www.desmos.com/calculator/vbwshyjoch That is the only point on the graph...